Ms. Patricia Sullivan plans to create a fund from her lottery winnings to meet three objectives. First, she wants to create a fund so that her mother can withdraw $20,000 per month for the remainder of her expected life of 20 years. Second, she wants to pay the down payment for her brother to buy a house upon graduation from college four years from now. She expects that he will need $100,000 for down payment at that time. Finally, she wants to retire after 15 years and be able to withdraw $30,000 per month starting a month from her retirement. She expects to live for 30 years after retirement. All monies earn 8 percent compounded monthly and all cash flows occur at the end of the relevant period.


How much money does she need to invest today to meet her first objective?

How much money does she need to invest today to meet all three objectives?
Answer #1
Objective 1
Excel formula Present value
=PV(8%/12,12*20,-20000,0,0) $ 23,91,085.83
Objective 2
Excel formula Present value
=PV(8%/12,4*12,0,-100000,0) $       72,692.06
Objective 3
Excel formula Present value
=PV(8%/12,12*30,-30000,0,0) ₹ 40,88,504.82
Excel formula Present value
=PV(8%/12,15*12,0,-4088504.85,0) $ 12,36,347.72
How much money does she need to invest today to meet her first objective?
Answer : $ 2,391,085.83
How much money does she need to invest today to meet all three objectives?
Answer : $ 2,391,085.83 + 72,692.06 + 1,236,347.72 = $3,700,125.61